# ResearStudio: A Human-Intervenable Framework for Building Controllable Deep-Research Agents

### v1

@@ Abstract
Current deep-research agents run in a “fire-and-forget” mode: once started, they give users no way to fix errors or add expert knowledge during execution. We present ResearStudio, the first open-source framework that places real-time human control at its core. The system follows a Collaborative Workshop design. A hierarchical Planner–Executor writes every step to a live “plan-as-document,” and a fast communication layer streams each action, file change, and tool call to a web interface. At any moment, the user can pause the run, edit the plan or code, run custom commands, and resume – switching smoothly between AI-led, human-assisted and human-led, AI-assisted modes. In fully autonomous mode, ResearStudio achieves state-of-the-art results on the GAIA benchmark, surpassing systems like OpenAI’s DeepResearch and Manus. These results show that strong automated performance and fine-grained human control can coexist. The full code, protocol, and evaluation scripts are available at https://github.com/ResearAI/ResearStudio. We will continue to update the repository to encourage further work on safe and controllable research agents. [^1] Our live demo is publicly accessible at http://ai-researcher.net:3000/.. [^2] We support the development of DeepScientist, which can be accessed at https://github.com/ResearAI/DeepScientist.
@@ 1 Introduction
The advent of Large Language Models (LLMs) (Ouyang et al., 2022; Achiam et al., 2023; Yang et al., 2025) has catalyzed a new era in artificial intelligence, providing powerful engines for reasoning (Weng et al., 2022; Besta et al., 2024), comprehension (Rein et al., 2024; Weng et al., 2024), and generation (Zhu et al., 2024; Li et al., 2024). This has naturally led to the development of LLM-based autonomous agents (Roy et al., 2024; Huang et al., 2025), which leverage these models as a cognitive core to tackle complex (Yao et al., 2023). These long-horizon tasks were previously intractable (Zhang et al., 2025a). Most recently, a new class of advanced autonomous systems, termed Deep Research (DR) agents, has emerged, exemplified by industry-leading solutions such as OpenAI DR (OpenAI, 2025), Gemini DR (Team et al., 2023), and Memento (Zhou et al., 2025).
Yet prevailing agent frameworks (Chen et al., 2024; Zhu et al., 2025; Zheng et al., 2025; OpenAI, 2025) offer only a rigid, one-directional pipeline: once a task is issued, the user is reduced to a passive observer. When the agent misinterprets goals or pursues flawed strategies, there is no channel for timely human intervention, leading to errors, wasted compute, and diminished trust. Current Deep Research agents, therefore, fall short of the collaborative interface envisioned for AI.
@@ 2 Related Work
The development of autonomous agents has been accelerated by foundational LLM advancements in reasoning, such as Chain-of-Thought (Wei et al., 2022) and self-reflection (Shinn et al., 2023), and in action, through tool-use integration (Schick et al., 2023). Building on this, agent architectures have explored multi-agent coordination, as in AutoGen (Wu et al., ), LangGraph and MetaGPT (Hong et al., ), or hierarchical task decomposition, such as in AgentOrchestra (Zhang et al., 2025b) and OWL (Hu et al., 2025). While these frameworks significantly advance agent-to-agent and agent-to-environment interactions, they largely overlook the paradigm of direct, real-time human-agent collaboration. ResearStudio, in contrast, applies these foundational reasoning and tool-use capabilities within an architecture designed specifically to place the human user at the center of the workflow.
@@ 3.3 Backend Models
ResearStudio (Pass@1) 77.36 69.77 61.54 70.91
Table 2: Performance comparison of our agent against baseline methods ODR-smolagents (Roucher et al., 2025), AutoAgent (Chen et al., 2024), OWL (Hu et al., 2025), A-World (at Ant Group, 2025), and OpenAI-DeepResearch (OpenAI, 2025) on GAIA benchmark (Mialon et al., 2023).
Agent Level-1 Level-2 Level-3 Average
OWL (Hu et al., 2025) 75.27 61.01 32.65 60.80
A-World (at Ant Group, 2025) 80.65 64.78 24.49 63.12
ResearStudio (Ours) 84.95 72.33 59.18 74.09
Table 3: Performance comparison of our agent against open-source frameworks on the test set of the GAIA benchmark.
@@ 4 Experiments
Following Memento (Zhou et al., 2025), we conducted a rigorous evaluation on the GAIA benchmark (Mialon et al., 2023), a standard for testing general-purpose agentic systems on complex reasoning and multi-step tool use. All experiments were performed in a fully autonomous mode, with no human intervention from task initiation to completion, thereby assessing the core performance of our architecture. For our agent’s configuration, the Planner is powered by gpt-4.1, while the Executor utilizes o4-mini for its tactical operations. To handle multimodal tasks, the framework is equipped with specialized models: gpt-4o for image-related tasks, gemini-2.5-pro for video processing, and Assembly AI for audio tasks. For evaluation, we report the Exact Match (EM) metric, where a prediction is considered correct only if it exactly matches the reference answer after normalizing for case, punctuation, and articles. The EM score is defined as the percentage of answers that achieve a perfect match.
@@
Our experimental results demonstrate that ResearStudio achieves state-of-the-art performance on the GAIA benchmark across both validation and test sets. As shown in Table 2, on the GAIA validation set, our framework achieves a leading average score of 70.91%, outperforming other established agents such as A-World (69.70%) and OpenAI-DeepResearch (67.36%). Notably, ResearStudio shows exceptional capability on the more complex tasks, achieving the highest scores on both Level-2 (69.77%) and the highly challenging Level-3 (61.54%) tasks.
To further validate these findings on unseen data, we evaluated our agent on the GAIA test set. The results, presented in Table 3, solidify ResearStudio’s superior performance. Our agent achieves an overall average score of 74.09%, surpassing all listed baselines across every difficulty level, with scores of 84.95% on Level-1, 72.33% on Level-2, and 59.18% on Level-3. These robust results, obtained in a fully autonomous setting, validate the efficacy of our Planner-Executor architecture and modular tool integration, proving that a framework designed for human-collaboration can also deliver superior performance on its own.
@@ 6 Conclusion
These limitations define clear avenues for our future work. To mitigate cognitive load and broaden the framework’s accessibility, we plan to develop semi-autonomous intervention mechanisms, such as AI-powered alerts that flag potential errors or logical inconsistencies for human review. To empirically validate the "Collaborative Workshop" paradigm, a critical next step is to conduct formal Human-Computer Interaction (HCI) studies. These studies will measure key metrics—such as task completion times with and without intervention, error correction rates, and user satisfaction scores—to provide direct evidence of collaborative utility. Lastly, to bolster the system’s robustness, we will undertake dedicated red-teaming efforts to assess its resilience against threats like prompt injection, data exfiltration, and the generation of harmful content, using the findings to engineer more sophisticated, dynamic safety protocols.

### v2

@@ Abstract
Current deep-research agents run in a “fire-and-forget” mode: once started, they give users no way to fix errors or add expert knowledge during execution. We present ResearStudio, the first open-source framework that places real-time human control at its core. The system follows a Collaborative Workshop design. A hierarchical Planner–Executor writes every step to a live “plan-as-document,” and a fast communication layer streams each action, file change, and tool call to a web interface. At any moment, the user can pause the run, edit the plan or code, run custom commands, and resume – switching smoothly between AI-led, human-assisted and human-led, AI-assisted modes. In fully autonomous mode, ResearStudio achieves state-of-the-art results on the GAIA benchmark, surpassing systems like OpenAI’s DeepResearch and Manus. These results show that strong automated performance and fine-grained human control can coexist. The full code, protocol, and evaluation scripts are available at https://github.com/ResearAI/ResearStudio. We will continue to update the repository to encourage further work on safe and controllable research agents. [^1] Our live demo is publicly accessible at http://ai-researcher.net:3000/. [^2] We support the development of DeepScientist, which can be accessed at https://github.com/ResearAI/DeepScientist.
@@ 1 Introduction
The advent of Large Language Models (LLMs) (ouyang2022training; achiam2023gpt; yang2025qwen3) has catalyzed a new era in artificial intelligence, providing powerful engines for reasoning (weng2022large; besta2024graph), comprehension (rein2024gpqa; wengmastering), and generation (zhu2024knowagent; li2024generation). This has naturally led to the development of LLM-based autonomous agents (roy2024exploring; huang2025deep), which leverage these models as a cognitive core to tackle complex (yao2023react). These long-horizon tasks were previously intractable (zhang2025darwin). Most recently, a new class of advanced autonomous systems, termed Deep Research (DR) agents, has emerged, exemplified by industry-leading solutions such as OpenAI DR (OpenAIDeepRS), Gemini DR (team2023gemini), and Memento (zhou2025mementofinetuningllmagents).
Yet prevailing agent frameworks (chen2024autoagents; zhu2025deepreview; zheng2025deepresearcher; OpenAIDeepRS) offer only a rigid, one-directional pipeline: once a task is issued, the user is reduced to a passive observer. When the agent misinterprets goals or pursues flawed strategies, there is no channel for timely human intervention, leading to errors, wasted compute, and diminished trust. Current Deep Research agents, therefore, fall short of the collaborative interface envisioned for AI.
@@ 2 Related Work
The development of autonomous agents has been accelerated by foundational LLM advancements in reasoning, such as Chain-of-Thought (wei2022chain) and self-reflection (shinn2023reflexion), and in action, through tool-use integration (schick2023toolformer). Building on this, agent architectures have explored multi-agent coordination, as in AutoGen (wuautogen), LangGraph and MetaGPT (hongmetagpt), or hierarchical task decomposition, such as in AgentOrchestra (zhang2025agentorchestra) and OWL (hu2025owl). While these frameworks significantly advance agent-to-agent and agent-to-environment interactions, they largely overlook the paradigm of direct, real-time human-agent collaboration. ResearStudio, in contrast, applies these foundational reasoning and tool-use capabilities within an architecture designed specifically to place the human user at the center of the workflow.
@@ 3.3 Backend Models
ResearStudio (zhou2025mementofinetuningllmagents) 75.47 70.93 53.85 69.70
Table 2: Performance comparison of our agent against baseline methods ODR-smolagents (roucher2025smolagents), AutoAgent (chen2024autoagents), OWL (hu2025owl), A-World (award), and OpenAI-DeepResearch (OpenAIDeepRS) on GAIA benchmark (mialon2023gaia).
@@ 4 Experiments
Following Memento (zhou2025mementofinetuningllmagents), we conducted a rigorous evaluation on the GAIA benchmark (mialon2023gaia), a standard for testing general-purpose agentic systems on complex reasoning and multi-step tool use. All experiments were performed in a fully autonomous mode, with no human intervention from task initiation to completion, thereby assessing the core performance of our architecture. For our agent’s configuration, the Planner is powered by gpt-4.1, while the Executor utilizes o4-mini for its tactical operations. To handle multimodal tasks, the framework is equipped with specialized models: gpt-4o for image-related tasks, gemini-2.5-pro for video processing, and Assembly AI for audio tasks. For evaluation, we report the Exact Match (EM) metric, where a prediction is considered correct only if it exactly matches the reference answer after normalizing for case, punctuation, and articles. The EM score is defined as the percentage of answers that achieve a perfect match.
@@
Our experimental results demonstrate that ResearStudio achieves state-of-the-art performance on the GAIA benchmark across both validation and test sets. As shown in Table 2, on the GAIA validation set, our framework achieves a leading average score of 70.91%, outperforming other established agents such as A-World (69.70%) and OpenAI-DeepResearch (67.36%). Notably, ResearStudio demonstrates exceptional capability in more complex tasks, achieving the highest scores on both Level 2 (69.77%) and the highly challenging Level 3 (61.54%) tasks. These robust results, obtained in a fully autonomous setting, validate the efficacy of our Planner-Executor architecture and modular tool integration, demonstrating that a framework designed for human collaboration can also deliver superior performance on its own.
@@ 6 Conclusion
These limitations define clear avenues for our future work. To mitigate cognitive load and broaden the framework’s accessibility, we plan to develop semi-autonomous intervention mechanisms, such as AI-powered alerts that flag potential errors or logical inconsistencies for human review. To empirically validate the “Collaborative Workshop” paradigm, a critical next step is to conduct formal Human-Computer Interaction (HCI) studies. These studies will measure key metrics—such as task completion times with and without intervention, error correction rates, and user satisfaction scores – to provide direct evidence of collaborative utility.
Figure 4: A successful execution trace of a GAIA Level-3 computational puzzle within the ResearStudio UI, showcasing the agent’s ability to write and execute code to find a correct solution.